# GENESIS-RL: GEnerating Natural Edge-cases with Systematic Integration of Safety considerations and Reinforcement Learning

### v1

@@ Abstract
In the rapidly evolving field of autonomous systems, the safety and reliability of the system components are fundamental requirements. These components are often vulnerable to complex and unforeseen environments, making natural edge-case generation essential for enhancing system resilience. This paper presents GENESIS-RL, a novel framework that leverages system-level safety considerations and reinforcement learning techniques to systematically generate naturalistic edge cases. By simulating challenging conditions that mimic the real-world situations, our framework aims to rigorously test entire system’s safety and reliability. Although demonstrated within the autonomous driving application, our methodology is adaptable across diverse autonomous systems. Our experimental validation, conducted on high-fidelity simulator underscores the overall effectiveness of this framework.
@@ I INTRODUCTION
Scenario-based testing is one of the key approaches for the validation of autonomous systems, especially those that incorporate learning-enabled components, that are known to be susceptible to rare, unexpected (potentially out-of-distribution) scenarios [1]. This testing approach is vital not only for ensuring the safety and reliability of these systems but also for enabling them to identify and rectify potential failures in diverse, unforeseen situations. In this context, identifying and preparing for challenging or edge-case scenarios becomes critical. Synthesizing realistic edge-case samples and incorporating these into the training process [2, 3] can significantly enhance the resilience of learning-enabled modules against adversarial conditions. By exposing the learning modules to these pessimistic samples, systems gain the opportunity to learn from challenging data and better generalize across a spectrum of real-world conditions. However, given the vast amount of possible scenarios, manual creation of every scenario is infeasible, making automated edge-case generation crucial for scalability and effectiveness [4].
In this regard, traditional adversarial attacks on machine learning models explore the vulnerability of the models by injecting imperceptible noise into the inputs [5, 6]. These input perturbation methods, while effective in degrading the model performance, typically generates unnatural and unrealistic samples, diverging from genuine real-world scenarios. Furthermore, these adversarial approaches usually target specific components of an autonomous system rather than assessing the system as a whole. This narrow focus can overlook the holistic behavior of the system, where, for instance, a failure in the perception module might be compensated by the system’s control mechanisms, thus not leading to a failure at the system level. On the other hand, if the control system is not able to compensate, a relatively small error in the perception module may lead to a catastrophic system-level failure. This highlights the limitation of focusing solely on component-level vulnerabilities without considering the integrated operation of the entire system.
@@
In this paper, we aim to alleviate these challenges, by performing edge-case generation with system-level safety objectives while maintaining the naturalness of the generated scenarios. We employ the rulebook formalism [17] to precisely specify system-level safety objectives and leverage the capabilities of Reinforcement Learning (RL) to guide the generation of scenarios that not only challenge the system across all its components but also resemble real-world conditions closely. By focusing on the end-to-end vulnerability of autonomous systems, our approach aims to generate scenarios where the system fails to adhere to rulebook safety rules, thereby identifying potential systemic failures. Also, our proposed framework ensures that the generated scenarios are not only challenging but also devoid of unrealistic artifacts (via use of high-fidelity simulators), offering a more effective and comprehensive approach to testing and validating the safety and reliability of autonomous systems. Lastly, while this study primarily showcases our approach within the autonomous driving (AD) domain, our approach is designed to be applicable to other learning-enabled autonomous systems.
@@
  • We provide extensive experimental validation of our framework in the context of the autonomous driving problems using the CARLA simulator [18]. We also demonstrate that a pre-trained RL policy can generate edge-cases for new scenarios with minimal to no training steps, thus accelerating the process of assessment and verification of learning-enabled autonomous systems.
@@ II Related Works
Recent research has explored diverse approaches to generating edge cases, each with unique contributions and limitations:
  • Cost Function-Based Identification: Efforts using cost functions to pinpoint high-risk scenarios have shown potential but often neglect critical factors like occlusions and unpredictable trajectories [19, 20, 21].
  • Perception-based techniques: Methods such as constant norm-based perturbation target the system’s perception capabilities but may not address the system’s overall performance comprehensively [22, 2, 7, 23].
  • Computer-vision based models: While innovative, methods that extract and recreate accidents from videos face challenges in accurately replicating real-world complexity [24].
  • Edge case generation toolkits: Some edge-case generation software toolkits, like VerifAI [25], are capable of analysis, falsification and data augmentation for systems utilizing ML architectures. It leverages an “abstract feature space” of higher-level information compared to the low-level “concrete feature space” of image pixels, to search for rule violation scenarios in a given environment.
@@ III Background
In this section, we introduce some key terminologies and foundational concepts, setting the stage for understanding our approach.
In this work, a system refers to the entity that is being evaluated for its ability to navigate and perform tasks within variable conditions. It could be an autonomous vehicle, robotic platform, or any computational model. The world denotes the simulated surroundings in which the system operates, a construct designed to emulate real-world dynamics where every aspect can have an effect on the system’s behavior. Lastly, an actor is an entity other than the system that also lives in the world.
As an example, in an autonomous driving context, the system here could be the ego vehicle, and the world is where the ego vehicle is situated. Other entities such as other vehicles and pedestrians on the street are actors. Together with crucial factors such as weather and road conditions (including road markings and traffic signs) that are not part of the system but could affect the system’s behavior, they are all parts of the world.
@@ III-A Deep Reinforcement Learning
Deep Reinforcement Learning (DRL) [26] is an extension of RL that harnesses the representational power of deep neural networks. At its core, DRL adheres to the Markov Decision Process (MDP) framework, mathematically formulated as a 4-tuple $(S,A,T,R)$, where:
  • $S$ represents the state space, comprising all conceivable states.
  • $A$ denotes the action space, encompassing all actions available to the agent.
  • $T:S\times A\times S^{\prime}\rightarrow[0,1]$ is the transition function, indicating the probability of transitioning from one state $S$ to another state $S^{\prime}$ given an action $A$.
@@
In a typical DRL setup, there is a DRL agent and the environment. The DRL agent is the entity that we hope to train for, whereas the environment is the setting or domain within which the DRL agent operates and makes decisions, which encompasses all aspects mentioned above in the MDP framework, including the state space $S$, action space $A$, the transition function $T$ and rewards $R$.
Under this MDP framework, the agent’s decision-making strategy at any time $t$ is governed by a policy $\pi$, which maps the current perceived state $s_{t}$ to an action $a_{t}$. In DRL, this policy is represented with neural networks, denoted as $\pi_{\theta}$, where $\theta$ represents the neural network’s trainable parameters. This configuration enables the agent to dynamically refine its strategy by updating $\theta$, thus improving its performance and adaptability in navigating the environment.
The objective of DRL is to discover an optimal policy $\pi_{\theta}^{*}$ that guides the agent to maximize the expected return along a trajectory $\tau$, which is a sequence of states and actions $(s_{0},a_{0},s_{1},a_{1},...,s_{T},a_{T})$. The expected return is calculated as $J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t},s_{t+1})]$, where $\gamma$ is the discount factor and $T$ the length of the trajectory. The concept of episodes emerge naturally from this setup. An episode describes a complete trajectory from an initial state to a terminal state [27].
@@ III-B Rulebook
We will use the rulebook formalism [17] to precisely describe the correct behavior of the system. A rulebook consists of a set $R$ of rules, each is evaluated over realizations. A realization is defined as a sequence of states of the system and all the other actors and features in the world. Given a set $\Sigma$ of realizations, a rule is defined as a function $r:\Sigma\to\mathbb{R}_{\geq 0}$ that measures the degree of violation of its argument. If $r(x)<r(y)$, then the realization $y$ violates the rule $r$ to a greater extent than does $x$. In particular, $r(x)=0$ indicates that a realization $x$ is fully compliant with the rule. Note that the definition of the violation metric might be analytical, “from first principles”, or be the result of a learning process.
In this work, we utilize the rulebook to calculate the rewards. A higher violation score leads to increased rewards, encouraging the agent to explore scenarios that challenge the system’s safety protocols and resilience, thereby generating critical edge cases.
@@ IV Methodology
In this section, we present how we formulate natural edge-case generation for an autonomous system, as a DRL problem along with our proposed algorithmic framework.
At a high level, GENESIS-RL utilizes DRL to dynamically explore and manipulate the conditions of a simulated world, aimed at generating challenging yet naturalistic edge-cases for a system. To achieve this, we have parameterized the world with parametric knobs—adjustable settings that control various aspects of the simulation, which in the case of autonomous driving, could include dynamic weather patterns, object placements, traffic flow, and so on. By adjusting these knobs, the DRL agent is provided with the capability to systematically probe and alter the simulated world, effectively simulating different edge cases that the system under test might encounter.
@@ IV-A1 State space
The state space encompasses all conceivable states $s_{t}$, which include permutations of parametric knobs, system’s behaviors, other actors, and features of the world. This state representation captures the dynamics of the world and the DRL agent’s action inputs, and is conveyed through information obtained by the system.
@@ IV-A3 Reward
The reward mechanism is designed to motivate the DRL agent to discover significant edge cases. It comprises two components: the learning module loss $r_{m}$ and the violation score $r_{v}$ derived from the rulebook. The learning module loss is the loss experienced by the learning-enabled module within the system, which acts a direct reward to the agent, where a lower loss indicates better performance of the module at performing its designated task. The violation score is an indirect reward provided to the agent due to the imperfection of the learning-enabled modules. For example, in the context of autonomous driving, the rulebook evaluates the ego vehicle’s trajectory against a set of predefined rules, penalizing actions that lead to unsafe scenarios. The total reward $r_{t}$ at time step $t$ is calculated as a combination of these two elements.
@@ IV-B GENESIS-RL Framework
To effectively implement our DRL formulation, we designed a framework consisting of the following components: the DRL agent, the initial scene generator, the simulator, the system, and the reward calculator. The latter four together forms an environment for the DRL agent, facilitating continuous learning of the DRL agent through dynamic interaction.
@@ IV-B1 DRL agent
The DRL agent is the decision-making core. At each time step $t$, it obtains the current state $s_{t}$ of the environment and executes an action $a_{t}$. The environment then responds to this action by evolving to a new state based on the updated parametric knobs of the simulated world and issues a scalar reward $r_{t}$ to the agent as a feedback.
@@ IV-B2 Initial scene generator
The initial scene generator is responsible for creating a distribution of the initial scenes (a configuration of physical objects, system and actors) and sampling from them in the simulated world. It dictates the initial conditions the system will encounter, therefore determining the initial scene observed by the DRL agent.
@@ IV-B3 Simulator
The simulator provides a realistic and interactive backdrop where the DRL agent’s actions and the system’s outputs are executed and new frames are updated, reflecting the changes in real time.
@@ IV-B4 System
The system is the entity being evaluated within variable conditions, as defined in the background section.
@@ IV-B5 Reward calculator
The reward calculator calculates the reward $r_{t}$ for time step $t$, as defined in the previous section.
@@ IV-C Training the DRL Agent
Now putting things together, a single step the training looks like as follows (See Fig. 1): at each time step $t$, the DRL agent receives a state $s_{t}$ from the simulator in the environment and executes an action $a_{t}$ on the simulator. The simulator reflects the changes based on the updated parametric knobs in the simulated world and the changes are subsequently captured by the system through its sensors. The system then generates its control signals based on its inputs, which leads to a system trajectory update. The updated trajectory is then evaluated by the rulebook for violation score calculation, and is sent back to the DRL agent combined with the learning module loss.
@@
Figure 1: Architectural overview of the proposed framework. At each step $t$, the DRL agent observes a state $s_{t}$ of the environment (1) and executes an action $a_{t}$ (2). The simulator then updates the simulated world accordingly and creates an updated frame. The updated frame (3) is then processed by the system to generate vehicle control signals (4). The control signals are subsequently applied to the simulated world to update the vehicle trajectory (5). The reward calculator evaluates the performance by comparing the ego vehicle’s trajectory against the rulebook and also computes the learning module loss, issuing a scalar reward $r_{t}$ (6) that guides the DRL agent’s learning process.
@@ V Experiments
In this paper, we explore the weather conditions that can lead to natural edge cases for autonomous driving. Hence, we grant the DRL agent exclusive control over the weather conditions in the simulated world. The system we evaluate is an ego vehicle, tasked to navigate through the simulated world based on sensor feedback. In subsequent sections, we detail the operationalization of the GENESIS-RL framework’s components, starting with the DRL agent and encompassing the environment, then the training and testing setups and evaluation metrics.
@@ V-A DRL Agent
We have chosen the Proximal Policy Optimization (PPO) algorithm [28] as the DRL agent for its stabilized training capabilities and its proficiency in handling continuous state and action spaces. We adopted the implementation of Stable-Baselines3 [29]. The state and action spaces for the agent is as follows:
@@
  • Action space: A 5-tuple that represents the delta in value for each parametric knob. The parametric knobs are: the fog density, the precipitation density, the precipitation deposit level, the sun altitude angle, and the sun azimuth angle. These actions are bounded between $[-1,1]$, and are scaled by a set of preset scalar factors to satisfy the perturbation limit, which is 5%. For example, the fog density ranges from 0 to 100, then the corresponding parametric knob would be scaled by a factor of 5 such that the change in fog density level between each step in an episode will not exceed 5%.
As mentioned earlier, the DRL agent is developing a policy $\pi_{\theta}$ parameterized by a neural network $\theta$. Here, the architecture of $\theta$ is a CNN, which takes the RGB three channel arrays $(640,480,3)$ as inputs, and outputs an 5-tuple action.

### v2

@@ Abstract
In the rapidly evolving field of autonomous vehicles, the safety and reliability of the system components are fundamental requirements. These components are often vulnerable to complex and unforeseen environments, making natural edge-case generation essential for enhancing system resilience. This paper presents GENESIS-RL, a novel framework that leverages system-level safety considerations and reinforcement learning techniques to systematically generate naturalistic edge cases. By simulating challenging conditions that mimic the real-world situations, our framework aims to rigorously test entire system’s safety and reliability. Our experimental validation, conducted on high-fidelity simulator underscores the overall effectiveness of this framework.
@@ I Introduction
Scenario-based testing is one of the key approaches for the validation of autonomous systems, especially those that incorporate learning-enabled components that are known to be susceptible to rare, unexpected (potentially out-of-distribution) scenarios [1]. This testing approach is vital not only for ensuring the safety and reliability of these systems but also for enabling them to identify and rectify potential failures in diverse, unforeseen situations. In this context, identifying and preparing for challenging or edge-case scenarios becomes critical. Synthesizing realistic edge-case samples and incorporating these into the training process [2, 3] can significantly enhance the resilience of learning-enabled modules against adversarial conditions. By exposing the learning modules to these pessimistic samples, systems gain the opportunity to learn from challenging data and better generalize across a spectrum of real-world conditions. However, given the vast amount of possible scenarios, manual creation of every scenario is infeasible, making automated edge-case generation crucial for scalability and effectiveness [4].
In this regard, traditional adversarial attacks on machine learning models explore the vulnerability of the models by injecting imperceptible noise into the inputs [5, 6]. These input perturbation methods, while effective in degrading the model performance, typically generates unnatural and unrealistic samples, diverging from genuine real-world scenarios. Furthermore, these approaches usually target specific components of an autonomous system rather than assessing the system as a whole. This narrow focus can overlook the holistic behavior of the system, where, for instance, a failure in the perception module might be compensated by the system’s control mechanisms, thus not leading to a failure at the system level. On the other hand, if the control system is not able to compensate, a relatively small error in the perception module may lead to a catastrophic system-level failure. This highlights the limitation of focusing solely on component-level vulnerabilities without considering the integrated operation of the entire system.
@@
In this paper, we aim to alleviate these challenges, by performing edge-case generation with system-level safety objectives while maintaining the naturalness of the generated scenarios. We employ the rulebook formalism [17] to precisely specify system-level safety objectives and leverage the capabilities of Reinforcement Learning (RL) to guide the generation of scenarios that not only challenge the system across all its components but also resemble real-world conditions closely. By focusing on the end-to-end vulnerability of autonomous systems, our approach aims to generate scenarios where the system fails to adhere to rulebook safety rules, thereby identifying potential systemic failures. Also, our proposed framework ensures that the generated scenarios are not only challenging but also devoid of unrealistic artifacts (via use of high-fidelity simulators), offering a more effective and comprehensive approach to testing and validating the safety and reliability of autonomous systems.
@@
  • We provide extensive experimental validation of our framework using the CARLA simulator [18]. We also demonstrate that a pre-trained RL policy can generate edge-cases for new scenarios with minimal to no training steps, thus accelerating the process of assessment and verification of learning-enabled autonomous systems.
@@ II Related Works
Recent research has explored diverse approaches to generating edge cases. Efforts using cost functions to pinpoint high-risk scenarios have shown potential yet often neglect critical factors like unpredictable trajectories [19, 20, 21]. Perception-based techniques, such as constant norm-based perturbation, target the system’s perception capabilities but may not address the system’s overall performance comprehensively [22, 2, 7, 23]. While innovative, methods that extract and recreate accidents from videos face challenges in accurately replicating real-world complexity [24].
Additionally, some edge-case generation software toolkits, like VerifAI [25], are capable of analysis, falsification, and data augmentation for systems utilizing ML architectures. These toolkits leverage an ”abstract feature space” of higher-level information compared to the low-level ”concrete feature space” of image pixels to search for rule violation scenarios in a given environment. Domain randomization effectively bridges the sim-to-real gap [26, 27], but it can lead to training an overly conservative policy depending on the range of randomization. System identification [28] offers a feasible solution by estimating environmental parameters through limited interaction with real-world scenarios.
Lastly, Bayesian optimization-based methods often generate challenging scenarios with limited diversity and insufficient complexity [2, 19]. These methods typically produce short scenario segments with limited environment interactions since they require a predefined parameter range [29]. Consequently, they limit the assessment of system performance and fail to capture realistic edge cases with diverse interactions.
@@ III Background
In this work, a system refers to the entity that is being evaluated for its ability to navigate and perform tasks within variable conditions. It could be an autonomous vehicle or any computational model. The world denotes the simulated surroundings in which the system operates, a construct designed to emulate real-world dynamics where every aspect can have an effect on the system’s behavior. Lastly, an actor is an entity other than the system that also lives in the world.
As an example, in an autonomous driving context, the system could be the ego vehicle, and the world is where the ego vehicle is situated. Other entities, such as other vehicles and pedestrians on the street, are actors. Together with crucial factors such as weather and road conditions (including road markings and traffic signs) that are not part of the system but could affect the system’s behavior, they are all parts of the world.
@@ III-A Deep Reinforcement Learning
Deep Reinforcement Learning (DRL) [30] is an extension of RL that harnesses the representational power of deep neural networks. At its core, DRL adheres to the Markov Decision Process (MDP) framework, mathematically formulated as a 4-tuple $(S,A,P,R)$, where:
  • $S$ represents the state space, comprising all conceivable states $s_{t}$ at a given time $t$.
  • $A$ denotes the action space, encompassing all actions $a_{t}$ available to the agent at time $t$.
  • $P$ the transition probability function, indicating the probability of transitioning from one state $s_{t}$ to another state $s_{t+1}$ given an action $a_{t}$.
@@
In a typical DRL setup, the DRL agent is the entity that we hope to train, whereas the environment is the setting or domain wherein the DRL agent operates and makes decisions, which encompasses all aspects mentioned above in the MDP framework, including the state space $S$, action space $A$, the transition probability function $P$ and rewards $R$.
Under this MDP framework, the agent’s decision-making strategy at any time $t$ is governed by a policy $\pi$, which maps the current state $s_{t}$ to an action $a_{t}$. In DRL, this policy is represented with neural networks, denoted as $\pi_{\theta}$, where $\theta$ represents the neural network’s trainable parameters. This configuration enables the agent to dynamically refine its strategy by updating $\theta$, thus improving its performance and adaptability in navigating the environment.
The objective of DRL is to discover an optimal policy $\pi_{\theta}^{*}$ that guides the agent to maximize the expected return along a trajectory $\tau$, which is a sequence of states and actions $(s_{0},a_{0},s_{1},a_{1},...,s_{T},a_{T})$. The expected return is calculated as $J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t},s_{t+1})]$, where $\gamma$ is the discount factor and $T$ the length of the trajectory. The concept of episodes emerges naturally from this setup. An episode describes a complete trajectory from an initial state to a terminal state [31].
@@ III-B Rulebook
We will use the rulebook formalism [17] to precisely describe the correct behavior of the system. A rulebook consists of a set $\Lambda$ of rules; each is evaluated over realizations. A realization is defined as a sequence of states of the system and all the other actors and features in the world. Given a set $\Sigma$ of realizations, a rule is defined as a function $\lambda:\Sigma\to\mathbb{R}_{\geq 0}$ that measures the degree of violation of its argument. If $\lambda(x)<\lambda(y)$, then the realization $y$ violates the rule $\lambda$ to a greater extent than does $x$. In particular, $\lambda(x)=0$ indicates that a realization $x$ is fully compliant with the rule. Note that the definition of the violation metric might be analytical, “from first principles”, or be the result of a learning process.
In this work, we utilize the rulebook to calculate the rewards. A higher violation score leads to increased rewards, encouraging the agent to explore scenarios that challenge the system’s safety protocols and resilience, thereby generating critical edge cases. For a detailed description of the rules used for reward calculation, please refer to the Reward calculator subsection in the Experiments section.
@@ IV Methodology
At a high level, GENESIS-RL utilizes DRL to dynamically explore and manipulate the conditions of a simulated world, aimed at generating challenging yet naturalistic edge-cases for a system. To achieve this, we parameterized the world with parametric knobs—adjustable settings that control various aspects of the simulation, which in the case of autonomous driving, could include dynamic weather patterns, object placements, traffic flow, and so on. By adjusting these knobs, the DRL agent is provided with the capability to systematically probe and alter the simulated world, effectively simulating different edge cases that the system under test might encounter.
@@ IV-A1 State space
The state space encompasses all conceivable states $s_{t}$, including permutations of parametric knobs, the system’s behaviors, other actors, and features of the world. This state representation captures the dynamics of the world and the DRL agent’s action inputs, and is conveyed through information obtained by the system.
@@ IV-A3 Reward
The reward mechanism is designed to motivate the DRL agent to discover edge cases. It comprises two components: the learning module loss $r_{m}$ and the violation score $r_{v}$ derived from the rulebook. The learning module loss is the loss experienced by the learning-enabled module within the system, which acts as a direct reward to the agent, where a lower loss indicates better performance of the module at performing its designated task. The violation score is an indirect reward provided to the agent due to the imperfection of the learning-enabled modules. For example, in autonomous driving, the rulebook evaluates the ego vehicle’s trajectory against a set of predefined rules, penalizing actions that lead to unsafe scenarios. The total reward $r_{t}$ at time step $t$ is calculated as a combination of these two elements.
@@ IV-B GENESIS-RL Framework
To implement our DRL formulation, we designed a framework consisting of the following components: the DRL agent, the initial scene generator, the simulator, the system, and the reward calculator. The latter four together form an environment for the DRL agent, facilitating continuous learning of the DRL agent through dynamic interaction.
@@ IV-B1 DRL agent
The DRL agent is the decision-making core. At each time step $t$, it obtains the current state $s_{t}$ of the environment and executes an action $a_{t}$. The environment then responds to this action by evolving to a new state based on the updated parametric knobs of the simulated world and issues a scalar reward $r_{t}$ to the agent as feedback.
@@ IV-B2 Initial scene generator
The initial scene generator is responsible for creating a distribution of the initial scenes (a configuration of physical objects, the system and actors) and sampling from them in the simulated world. It dictates the initial conditions the system will encounter, therefore determining the initial scene observed by the DRL agent.
@@ IV-B3 Simulator
The simulator provides a realistic and interactive backdrop where the DRL agent’s actions and the system’s outputs are executed and new frames are updated, reflecting the changes in real-time.
@@ IV-B4 System
As defined in the background section, The system is the entity being evaluated within variable conditions.
@@ IV-B5 Reward calculator
As defined in the previous section, the reward calculator calculates the reward $r_{t}$ for time step $t$.
@@ IV-C Training the DRL Agent
Putting things together, a single step of the training looks like as follows (See Fig. 1): at each time step $t$, the DRL agent receives a state $s_{t}$ from the simulator and executes an action $a_{t}$ on the simulator. The simulator reflects the changes based on the updated parametric knobs and the changes are subsequently captured by the system through its sensors. The system then generates control signals based on its inputs, which leads to a system trajectory update. The updated trajectory is then evaluated by the rulebook for violation score calculation and is sent back to the DRL agent combined with the learning module loss.
@@
Figure 1: Architectural overview of the proposed framework. At each step $t$, the DRL agent observes a state $s_{t}$ (1) and executes an action $a_{t}$ (2). The simulator then updates the simulated world accordingly and creates an updated frame. The updated frame (3) is then processed by the system to generate vehicle control signals (4). The control signals are then applied to the simulated world to update the vehicle trajectory (5). The reward calculator evaluates the performance by comparing the ego vehicle’s trajectory against the rulebook and also computes the learning module loss, issuing a scalar reward $r_{t}$ (6) that guides the DRL agent’s learning process.
@@ V Experiments
In this paper, we explore the weather conditions that can lead to natural edge cases for autonomous driving. Hence, we grant the DRL agent exclusive control over the weather conditions in the simulated world. The system we evaluate is the ego vehicle, tasked to navigate through the simulated world based on sensor feedback. In subsequent sections, we detail the operationalization of the GENESIS-RL framework’s components, starting with the DRL agent and encompassing the environment, then the training and testing setups and evaluation metrics.
Remark.
Despite only having weather parameters as available actions in this work, we can also include other factors such as the behavior of other actors into our framework by parameterizing their behavior in the simulation (such as the aggressiveness of the other actors using CARLA’s traffic manager, or the position and orientation of the actors using Scenic [32]).
@@ V-A DRL Agent
We have chosen the Proximal Policy Optimization (PPO) algorithm [33] as the DRL agent for its stabilized training capabilities and proficiency in handling continuous state and action spaces. We adopted the implementation of Stable-Baselines3 [34]. The state and action spaces are as follows:
@@
  • Action space: A 5-tuple that represents the delta in value for each parametric knob. The parametric knobs are: the fog density, the precipitation density, the precipitation deposit level, the sun altitude angle, and the sun azimuth angle. These actions are bounded between $[-1,1]$, and are scaled by a set of preset scalar factors to satisfy the perturbation limit, which is 5%. For example, if the fog density ranges from 0 to 100, then the corresponding parametric knob would be scaled by a factor of 5 such that the change in fog density level between each step in an episode will not exceed 5%.
As mentioned earlier, the DRL agent is developing a policy $\pi_{\theta}$ parameterized by a neural network $\theta$. Here, the architecture of $\theta$ is a CNN, which takes the RGB three channel arrays $(640,480,3)$ as inputs, and outputs a 5-tuple action.